# Linear Bandits with Partially Observable Features

### v2

@@ E.4 Proof of Theorem 4
$\widetilde{\mathbf{x}}_{a}^{\top}(\widehat{\boldsymbol{\mu}}_{t}^{R}-{\boldsymbol{\mu}}_{\star})=\widetilde{\mathbf{x}}_{a}^{\top}\widetilde{\mathbf{V}}_{t}^{-1}\left\{\left(\widetilde{\mathbf{V}}_{t}-\widetilde{\mathbf{A}}_{t}\right)\left(\check{\boldsymbol{\mu}}_{t}^{R}-{\boldsymbol{\mu}}_{\star}\right)+\sum_{\tau=1}^{t}\mathbb{I}(\mathcal{M}_{\tau})\sum_{a\in[K]}\frac{\mathbb{I}(\widetilde{a}_{\tau}=a)}{\phi_{a,\tau}}\widetilde{\mathbf{x}}_{a}\epsilon_{\tau}-{\boldsymbol{\mu}}_{\star}\right\}.$
@@
$\check{\boldsymbol{\mu}}_{t}^{R}-{\boldsymbol{\mu}}_{\star}=$ $\left(\sum_{\tau=1}^{t}\widetilde{\boldsymbol{x}}_{a_{\tau}}\widetilde{\boldsymbol{x}}_{a_{\tau}}^{\top}+p\mathbf{I}_{K}\right)^{-1}\left(\sum_{\tau=1}^{t}\widetilde{\boldsymbol{x}}_{a_{\tau}}\epsilon_{\tau}-p{\boldsymbol{\mu}}_{\star}\right)$
$=$ $\left(\sum_{\tau=1}^{t}\frac{1}{\phi_{a_{\tau},\tau}}\widetilde{\boldsymbol{x}}_{a_{\tau}}\widetilde{\boldsymbol{x}}_{a_{\tau}}^{\top}+\mathbf{I}_{K}\right)^{-1}\left(\sum_{\tau=1}^{t}\frac{1}{\phi_{a_{\tau},\tau}}\widetilde{\boldsymbol{x}}_{a_{\tau}}\epsilon_{\tau}-{\boldsymbol{\mu}}_{\star}\right)$
$=\left(\sum_{\tau=1}^{t}\sum_{a\in[K]}\frac{\mathbb{I}(\widetilde{a}_{\tau}=a)}{\phi_{a,\tau}}\widetilde{\boldsymbol{x}}_{a_{\tau}}\widetilde{\boldsymbol{x}}_{a_{\tau}}^{\top}+\mathbf{I}_{K}\right)^{-1}\left(\sum_{\tau=1}^{t}\frac{1}{p}\widetilde{\boldsymbol{x}}_{a_{\tau}}\epsilon_{\tau}-{\boldsymbol{\mu}}_{\star}\right),$
where the second equality holds because $\phi_{a_{\tau},\tau}=p$. Under the coupling event $\cap_{\tau=1}^{t}\mathcal{M}_{\tau}$,
$\sum_{\tau=1}^{t}\mathbb{I}(\mathcal{M}_{\tau})\sum_{a\in[K]}\frac{\mathbb{I}(\widetilde{a}_{\tau}=a)}{\phi_{a,t}}\widetilde{\mathbf{x}}_{a}\widetilde{\mathbf{x}}_{a}^{\top}+\mathbf{I}_{K}=$ $\sum_{\tau=1}^{t}\sum_{a\in[K]}\frac{\mathbb{I}(\widetilde{a}_{\tau}=a)}{\phi_{a,t}}\widetilde{\mathbf{x}}_{a}\widetilde{\mathbf{x}}_{a}^{\top}+\mathbf{I}_{K}$
::=:= $\mathbf{A}_{t},$
and
$\widetilde{\mathbf{x}}_{a}^{\top}(\widehat{\boldsymbol{\mu}}_{t}^{R}-{\boldsymbol{\mu}}_{\star})=$ $\widetilde{\mathbf{x}}_{a}^{\top}\mathbf{V}_{t}^{-1}\left\{\left(\mathbf{V}_{t}-\mathbf{A}_{t}\right)\mathbf{A}_{t}^{-1}\left(\sum_{\tau=1}^{t}\frac{1}{p}\widetilde{\boldsymbol{x}}_{a_{\tau}}\epsilon_{\tau}-{\boldsymbol{\mu}}_{\star}\right)+\sum_{\tau=1}^{t}\frac{\epsilon_{\tau}}{\phi_{a_{\tau},\tau}}\widetilde{\boldsymbol{x}}_{a_{\tau}}-{\boldsymbol{\mu}}_{\star}\right\}$
$=$ $\widetilde{\mathbf{x}}_{a}^{\top}\mathbf{V}_{t}^{-1}\left\{\left(\mathbf{V}_{t}-\mathbf{A}_{t}\right)\mathbf{A}_{t}^{-1}+\mathbf{I}_{K}\right\}\left(\sum_{\tau=1}^{t}\frac{1}{p}\widetilde{\boldsymbol{x}}_{a_{\tau}}\epsilon_{\tau}-{\boldsymbol{\mu}}_{\star}\right)$
$=$ $\widetilde{\mathbf{x}}_{a}^{\top}\mathbf{V}_{t}^{-1/2}\left(\mathbf{V}_{t}^{1/2}\mathbf{A}_{t}^{-1}\mathbf{V}_{t}^{1/2}\right)\mathbf{V}_{t}^{-1/2}\left(\sum_{\tau=1}^{t}\frac{1}{p}\widetilde{\boldsymbol{x}}_{a_{\tau}}\epsilon_{\tau}-{\boldsymbol{\mu}}_{\star}\right).$
@@
$\max_{a\in[K]}|{\widetilde{\mathbf{x}}_{a}^{\top}(\widehat{\boldsymbol{\mu}}_{t}^{R}-{\boldsymbol{\mu}}_{\star})}|\leq\max_{a\in[K]}\|\widetilde{\mathbf{x}}_{a}\|_{\mathbf{V}_{t}^{-1}}\|\mathbf{V}_{t}^{1/2}\mathbf{A}_{t}^{-1}\mathbf{V}_{t}^{1/2}\|_{2}\left\|\sum_{\tau=1}^{t}\frac{1}{p}\widetilde{\boldsymbol{x}}_{a_{\tau}}\epsilon_{\tau}-{\boldsymbol{\mu}}_{\star}\right\|_{\mathbf{V}_{t}^{-1}}.$
By Corollary which implies $\mathbf{I}_{K}-\mathbf{V}_{t}^{-1/2}\mathbf{A}_{t}\mathbf{V}_{t}^{-1/2}\preceq\epsilon\mathbf{I}_{K}.$ Rearraging the terms,
@@
Thus,
$\max_{a\in[K]}|{\widetilde{\mathbf{x}}_{a}^{\top}(\widehat{\boldsymbol{\mu}}_{t}^{R}-{\boldsymbol{\mu}}_{\star})}|$ $\leq\frac{\max_{a\in[K]}\|\widetilde{\mathbf{x}}_{a}\|_{\mathbf{V}_{t}^{-1}}}{1-\epsilon}\left\|\sum_{\tau=1}^{t}\frac{1}{p}\widetilde{\boldsymbol{x}}_{a_{\tau}}\epsilon_{\tau}-{\boldsymbol{\mu}}_{\star}\right\|_{\mathbf{V}_{t}^{-1}}$
$\leq\frac{\max_{a\in[K]}\|\widetilde{\mathbf{x}}_{a}\|_{\mathbf{V}_{t}^{-1}}}{1-\epsilon}\left(\frac{1}{p}\left\|\sum_{\tau=1}^{t}\widetilde{\boldsymbol{x}}_{a_{\tau}}\epsilon_{\tau}\right\|_{\mathbf{V}_{t}^{-1}}+\left\|{\boldsymbol{\mu}}_{\star}\right\|_{\mathbf{V}_{t}^{-1}}\right).$
Note that the matrix $\mathbf{V}_{t}$ is deterministic. By Lemma 9 in (Abbasi-yadkori et al., 2011), with probability at least $1-\delta,$
$\left\|\sum_{\tau=1}^{t}\widetilde{\boldsymbol{x}}_{a_{\tau}}\epsilon_{\tau}\right\|_{\mathbf{V}_{t}^{-1}}$ $\leq\left\|\sum_{\tau=1}^{t}\widetilde{\boldsymbol{x}}_{a_{\tau}}\epsilon_{\tau}\right\|_{\left(\sum_{\tau=1}^{t}\widetilde{\boldsymbol{x}}_{a_{\tau}}\widetilde{\boldsymbol{x}}_{a_{\tau}}^{\top}+\mathbf{I}_{K}\right)^{-1}}$
$\leq\sigma\sqrt{2\log\frac{\det(\sum_{\tau=1}^{t}\widetilde{\boldsymbol{x}}_{a_{\tau}}\widetilde{\boldsymbol{x}}_{a_{\tau}}^{\top}+\mathbf{I}_{K}){}^{1/2}}{\delta}}$
$\leq\sigma\sqrt{\log\frac{\det(\sum_{\tau=1}^{t}\widetilde{\boldsymbol{x}}_{a_{\tau}}\widetilde{\boldsymbol{x}}_{a_{\tau}}^{\top}+\mathbf{I}_{K})}{\delta}},$
for all $t\geq 1$. Because
$\det\left(\sum_{\tau=1}^{t}\widetilde{\boldsymbol{x}}_{a_{\tau}}\widetilde{\boldsymbol{x}}_{a_{\tau}}^{\top}+\mathbf{I}_{K}\right)$$\leq\left\{\frac{\text{Tr}\left(\sum_{\tau=1}^{t}\widetilde{\boldsymbol{x}}_{a_{\tau}}\widetilde{\boldsymbol{x}}_{a_{\tau}}^{\top}\right)+K}{K}\right\}^{K}$
$\leq\left\{\frac{t\max_{a\in[K]}\|\widetilde{\boldsymbol{x}}_{a_{\tau}}\|_{2}+K}{K}\right\}^{K}$
@@
where the last inequality holds by $\|\widetilde{\boldsymbol{x}}_{a_{\tau}}\|_{2}\leq\sqrt{K}\|\widetilde{\boldsymbol{x}}_{a_{\tau}}\|_{\infty}\leq K$. Thus,
$\left\|\sum_{\tau=1}^{t}\widetilde{\boldsymbol{x}}_{a_{\tau}}\epsilon_{\tau}\right\|_{\mathbf{V}_{t}^{-1}}\leq\sigma\sqrt{K\log\frac{t+1}{\delta}},$
which proves,
@@
Because $\|{\boldsymbol{\mu}}_{\star}\|_{\mathbf{V}_{t}^{-1}}\leq\|{\boldsymbol{\mu}}_{\star}\|_{2}\leq\sqrt{K}$, setting $\epsilon=1/2$ completes the proof. ∎
E.5 Proof of Theorem 5
Because the regret is bounded by $1$ and the number of rounds for the exploration phase is at most $|\mathcal{E}_{T}|\leq 32(1-p)^{-2}K^{2}\log({2dT^{2}}/{\delta}).$
@@ E.5 Proof of Theorem 5
$=$ $32(1-p)^{-2}K^{2}\log\frac{2dT^{2}}{\delta}+\sum_{t\in[T]\setminus\mathcal{E}_{T}}\left\{\mathbb{I}\left(a_{t}=\widehat{a}_{t}\right)\left(\mathbb{E}_{t-1}[y_{\star,t}]-\mathbb{E}_{t-1}[y_{a_{t},t}]\right)\right\}$
$+\sum_{t\in[T]\setminus\mathcal{E}_{T}}\left\{\mathbb{I}\left(a_{t}\neq\widehat{a}_{t}\right)\left(\mathbb{E}_{t-1}[y_{\star,t}]-\mathbb{E}_{t-1}[y_{a_{t},t}]\right)\right\}.$
On the event $\{a_{t}=\widehat{a}_{t}\}$,
$\mathbb{E}_{t-1}[y_{\star,t}]-\mathbb{E}_{t-1}[y_{a_{t},t}]=$ $\widetilde{\boldsymbol{x}}_{a_{\star}}^{\top}{\boldsymbol{\mu}}_{\star}-\widetilde{\boldsymbol{x}}_{\widehat{a}_{t}}^{\top}{\boldsymbol{\mu}}_{\star}$
$\leq 2\max_{a\in[K]}\left|\widetilde{\mathbf{x}}_{a}^{\top}\left({\boldsymbol{\mu}}_{\star}-\widehat{\boldsymbol{\mu}}_{t-1}^{R}\right)\right|+\widetilde{\boldsymbol{x}}_{a_{\star}}^{\top}\widehat{\boldsymbol{\mu}}_{t-1}^{R}-\widetilde{\boldsymbol{x}}_{\widehat{a}_{t}}^{\top}\widehat{\boldsymbol{\mu}}_{t-1}^{R}$
@@
$\leq\frac{4}{\sqrt{t}}\left(\frac{\sigma}{p}\sqrt{K\log\frac{2t^{2}}{\delta}}+\sqrt{K}\right),$
with probability at least $1-3\delta/t$, by Theorem 4. Summing over $t$ gives, (with probability at least $1-3\delta$),
$\sum_{t\in[T]\setminus\mathcal{E}_{T}}\left\{\mathbb{I}\left(a_{t}=\widehat{a}_{t}\right)\left(\mathbb{E}_{t-1}[y_{\star,t}]-\mathbb{E}_{t-1}[y_{a_{t},t}]\right)\right\}\leq 8\sqrt{KT}\left(\frac{\sigma}{p}\sqrt{\log\frac{2T^{2}}{\delta}}+1\right).$
By Lemma 5 and the Hoeffding bound, with probability at least $1-\delta$
$\sum_{t\in[T]\setminus\mathcal{E}_{T}}\left\{\mathbb{I}\left(a_{t}\neq\widehat{a}_{t}\right)\left(\mathbb{E}_{t-1}[y_{\star,t}]-\mathbb{E}_{t-1}[y_{a_{t},t}]\right)\right\}$
$\leq 2\left[\sum_{t\in[T]}\mathbb{I}\left(a_{t}\neq\widehat{a}_{t}\right)-\mathbb{P}\left(a_{t}\neq\widehat{a}_{t}\right)+\mathbb{P}\left(a_{t}\neq\widehat{a}_{t}\right)\right]$
$\leq 4\sqrt{2T\log\frac{2}{\delta}}+\frac{4\sqrt{T}}{K-1}+4\delta.$
Then the cumulative regret,
$\mathrm{Reg}(T)\leq\frac{32K^{2}}{(1-p)^{2}}\log\frac{2dT^{2}}{\delta}+4\sqrt{2T\log\frac{2}{\delta}}+\frac{4\sqrt{T}}{K-1}+4\delta+8\sqrt{KT}\left(\frac{\sigma}{p}\sqrt{\log\frac{t+1}{\delta}}+1\right),$

### v3

@@ E.4 Proof of Theorem˜4
$\widetilde{\mathbf{x}}_{a}^{\top}(\widehat{\boldsymbol{\mu}}_{t}^{R}-{\boldsymbol{\mu}}_{\star})=\widetilde{\mathbf{x}}_{a}^{\top}\widetilde{\mathbf{V}}_{t}^{-1}\left\{\left(\widetilde{\mathbf{V}}_{t}-\widetilde{\mathbf{A}}_{t}\right)\left(\check{\boldsymbol{\mu}}_{t}^{R}-{\boldsymbol{\mu}}_{\star}\right)+\sum_{\tau=1}^{t}\mathbb{I}(\mathcal{M}_{\tau})\sum_{a\in[K]}\frac{\mathbb{I}(\widetilde{a}_{\tau}=a)}{\phi_{a,\tau}}\widetilde{\mathbf{x}}_{a}\epsilon_{\tau}-{\boldsymbol{\mu}}_{\star}\right\}.$ (28)
@@
$\check{\boldsymbol{\mu}}_{t}^{R}-{\boldsymbol{\mu}}_{\star}$ $=\left(\sum_{\tau=1}^{t}\widetilde{\mathbf{x}}_{a_{\tau}}\widetilde{\mathbf{x}}_{a_{\tau}}^{\top}+p\mathbf{I}_{K}\right)^{-1}\left(\sum_{\tau=1}^{t}\widetilde{\mathbf{x}}_{a_{\tau}}\epsilon_{\tau}-p{\boldsymbol{\mu}}_{\star}\right)$
$=\left(\sum_{\tau=1}^{t}\frac{1}{\phi_{a_{\tau},\tau}}\widetilde{\mathbf{x}}_{a_{\tau}}\widetilde{\mathbf{x}}_{a_{\tau}}^{\top}+\mathbf{I}_{K}\right)^{-1}\left(\sum_{\tau=1}^{t}\frac{1}{\phi_{a_{\tau},\tau}}\widetilde{\mathbf{x}}_{a_{\tau}}\epsilon_{\tau}-{\boldsymbol{\mu}}_{\star}\right)$
$=\left(\sum_{\tau=1}^{t}\sum_{a\in[K]}\frac{\mathbb{I}(\widetilde{a}_{\tau}=a)}{\phi_{a,\tau}}\widetilde{\mathbf{x}}_{a_{\tau}}\widetilde{\mathbf{x}}_{a_{\tau}}^{\top}+\mathbf{I}_{K}\right)^{-1}\left(\sum_{\tau=1}^{t}\frac{1}{p}\widetilde{\mathbf{x}}_{a_{\tau}}\epsilon_{\tau}-{\boldsymbol{\mu}}_{\star}\right),$
where the second equality holds because $\phi_{a_{\tau},\tau}=p$ and the coupling event $\cap_{\tau=1}^{t}\mathcal{M}_{\tau}$. Thus,
$\check{\boldsymbol{\mu}}_{t}^{R}-{\boldsymbol{\mu}}_{\star}=\mathbf{A}_{t}^{-1}\left(\sum_{\tau=1}^{t}\frac{1}{p}\widetilde{\mathbf{x}}_{a_{\tau}}\epsilon_{\tau}-{\boldsymbol{\mu}}_{\star}\right).$
Plugging in (28),
$\widetilde{\mathbf{x}}_{a}^{\top}(\widehat{\boldsymbol{\mu}}_{t}^{R}-{\boldsymbol{\mu}}_{\star})=\widetilde{\mathbf{x}}_{a}^{\top}\widetilde{\mathbf{V}}_{t}^{-1}\left\{\left(\widetilde{\mathbf{V}}_{t}-\widetilde{\mathbf{A}}_{t}\right)\mathbf{A}_{t}^{-1}\left(\sum_{\tau=1}^{t}\frac{1}{p}\widetilde{\mathbf{x}}_{a_{\tau}}\epsilon_{\tau}-{\boldsymbol{\mu}}_{\star}\right)+\sum_{\tau=1}^{t}\mathbb{I}(\mathcal{M}_{\tau})\sum_{a\in[K]}\frac{\mathbb{I}(\widetilde{a}_{\tau}=a)}{\phi_{a,\tau}}\widetilde{\mathbf{x}}_{a}\epsilon_{\tau}-{\boldsymbol{\mu}}_{\star}\right\}$
Under the coupling event, $\widetilde{\mathbf{A}}_{t}=\mathbf{A}_{t}$, $\widetilde{\mathbf{V}}_{t}=\mathbf{V}_{t}$ and $\sum_{a\in[K]}\phi_{a,\tau}^{-1}\mathbb{I}(\widetilde{a}_{\tau}=a)\widetilde{\mathbf{x}}_{a}={\widetilde{\mathbf{x}}_{a_{\tau}}}/{p}$. It follows that
$\widetilde{\mathbf{x}}_{a}^{\top}(\widehat{\boldsymbol{\mu}}_{t}^{R}-{\boldsymbol{\mu}}_{\star})$ $=\widetilde{\mathbf{x}}_{a}^{\top}\mathbf{V}_{t}^{-1}\left\{\left(\mathbf{V}_{t}-\mathbf{A}_{t}\right)\mathbf{A}_{t}^{-1}+\mathbf{I}_{K}\right\}\left(\sum_{\tau=1}^{t}\frac{1}{p}\widetilde{\mathbf{x}}_{a_{\tau}}\epsilon_{\tau}-{\boldsymbol{\mu}}_{\star}\right)$
$=\widetilde{\mathbf{x}}_{a}^{\top}\mathbf{V}_{t}^{-1/2}\left(\mathbf{V}_{t}^{1/2}\mathbf{A}_{t}^{-1}\mathbf{V}_{t}^{1/2}\right)\mathbf{V}_{t}^{-1/2}\left(\sum_{\tau=1}^{t}\frac{1}{p}\widetilde{\mathbf{x}}_{a_{\tau}}\epsilon_{\tau}-{\boldsymbol{\mu}}_{\star}\right).$
@@
$\max_{a\in[K]}|{\widetilde{\mathbf{x}}_{a}^{\top}(\widehat{\boldsymbol{\mu}}_{t}^{R}-{\boldsymbol{\mu}}_{\star})}|\leq\max_{a\in[K]}\|\widetilde{\mathbf{x}}_{a}\|_{\mathbf{V}_{t}^{-1}}\|\mathbf{V}_{t}^{1/2}\mathbf{A}_{t}^{-1}\mathbf{V}_{t}^{1/2}\|_{2}\left\|\sum_{\tau=1}^{t}\frac{1}{p}\widetilde{\mathbf{x}}_{a_{\tau}}\epsilon_{\tau}-{\boldsymbol{\mu}}_{\star}\right\|_{\mathbf{V}_{t}^{-1}}.$ (29)
Corollary˜1 implies that, for $t\geq 8\epsilon^{-2}(1-p)^{-2}K^{2}\log(2Kt^{2}/\delta)$, with $\epsilon$ to be determined later,
$\mathbf{I}_{K}-\mathbf{V}_{t}^{-1/2}\mathbf{A}_{t}\mathbf{V}_{t}^{-1/2}\preceq\epsilon\mathbf{I}_{K},$
with probability at least $1-\delta$ for all $t\geq 1$. Rearranging terms gives
@@
Combining this with Eq.˜29,
$\max_{a\in[K]}|{\widetilde{\mathbf{x}}_{a}^{\top}(\widehat{\boldsymbol{\mu}}_{t}^{R}-{\boldsymbol{\mu}}_{\star})}|$ $\leq\frac{\max_{a\in[K]}\|\widetilde{\mathbf{x}}_{a}\|_{\mathbf{V}_{t}^{-1}}}{1-\epsilon}\left\|\sum_{\tau=1}^{t}\frac{1}{p}\widetilde{\mathbf{x}}_{a_{\tau}}\epsilon_{\tau}-{\boldsymbol{\mu}}_{\star}\right\|_{\mathbf{V}_{t}^{-1}}$
$\leq\frac{\max_{a\in[K]}\|\widetilde{\mathbf{x}}_{a}\|_{\mathbf{V}_{t}^{-1}}}{1-\epsilon}\left(\frac{1}{p}\left\|\sum_{\tau=1}^{t}\widetilde{\mathbf{x}}_{a_{\tau}}\epsilon_{\tau}\right\|_{\mathbf{V}_{t}^{-1}}+\left\|{\boldsymbol{\mu}}_{\star}\right\|_{\mathbf{V}_{t}^{-1}}\right),$
where the last inequality follows from the triangle inequality. Note that $\mathbf{V}_{t}$ is deterministic, both $\mathbf{V}_{t}$ and $\sum_{\tau=1}^{t}\widetilde{\mathbf{x}}_{a_{\tau}}\widetilde{\mathbf{x}}_{a_{\tau}}^{\top}+\mathbf{I}_{K}$ are positive definite, and $\mathbf{V}_{t}\succeq\sum_{\tau=1}^{t}\widetilde{\mathbf{x}}_{a_{\tau}}\widetilde{\mathbf{x}}_{a_{\tau}}^{\top}+\mathbf{I}_{K}$. Then, by Lemma 9 of (Abbasi-yadkori et al., 2011), with probability at least $1-\delta$,
$\left\|\sum_{\tau=1}^{t}\widetilde{\mathbf{x}}_{a_{\tau}}\epsilon_{\tau}\right\|_{\mathbf{V}_{t}^{-1}}$ $\leq\left\|\sum_{\tau=1}^{t}\widetilde{\mathbf{x}}_{a_{\tau}}\epsilon_{\tau}\right\|_{\left(\sum_{\tau=1}^{t}\widetilde{\mathbf{x}}_{a_{\tau}}\widetilde{\mathbf{x}}_{a_{\tau}}^{\top}+\mathbf{I}_{K}\right)^{-1}}$
$\leq\sigma\sqrt{2\log\frac{\det(\sum_{\tau=1}^{t}\widetilde{\mathbf{x}}_{a_{\tau}}\widetilde{\mathbf{x}}_{a_{\tau}}^{\top}+\mathbf{I}_{K}){}^{1/2}}{\delta}}$
$\leq\sigma\sqrt{\log\frac{\det(\sum_{\tau=1}^{t}\widetilde{\mathbf{x}}_{a_{\tau}}\widetilde{\mathbf{x}}_{a_{\tau}}^{\top}+\mathbf{I}_{K})}{\delta}}$
Applying Lemma 10 of Abbasi-yadkori et al. (2011) yields
$\det\left(\sum_{\tau=1}^{t}\widetilde{\mathbf{x}}_{a_{\tau}}\widetilde{\mathbf{x}}_{a_{\tau}}^{\top}+\mathbf{I}_{K}\right)$ $\leq\left\{\frac{\text{Tr}\left(\sum_{\tau=1}^{t}\widetilde{\mathbf{x}}_{a_{\tau}}\widetilde{\mathbf{x}}_{a_{\tau}}^{\top}\right)+K}{K}\right\}^{K}$
$\leq\left\{\frac{t\max_{a\in[K]}\|\widetilde{\mathbf{x}}_{a_{\tau}}\|_{2}+K}{K}\right\}^{K}$
@@
for all $t\geq 1$, where the last inequality holds by $\|\widetilde{\mathbf{x}}_{a_{\tau}}\|_{2}\leq\sqrt{K}\|\widetilde{\mathbf{x}}_{a_{\tau}}\|_{\infty}\leq K$. Hence,
$\left\|\sum_{\tau=1}^{t}\widetilde{\mathbf{x}}_{a_{\tau}}\epsilon_{\tau}\right\|_{\mathbf{V}_{t}^{-1}}\leq\sigma\sqrt{K\log\frac{t+1}{\delta}},$
which follows that,
@@
Since $\|{\boldsymbol{\mu}}_{\star}\|_{\mathbf{V}_{t}^{-1}}\leq\|{\boldsymbol{\mu}}_{\star}\|_{2}\leq\sqrt{K}$, choosing $\epsilon=1/2$ completes the proof. ∎
E.5 Proof of Theorem˜5
Similar to the proof of Theorem˜3 (Section˜E.3), the instantaneous regret is bounded above by $2$ for any $t\in[T]$, and the number of rounds for the exploration phase satisfies $|\mathcal{E}_{T}|\leq 32(1-p)^{-2}K^{2}\log({2dT^{2}}/{\delta})$. Consequently, the cumulative regret is bounded above as follows:
@@ E.5 Proof of Theorem˜5
$\begin{aligned} &=32(1-p)^{-2}K^{2}\log\frac{2dT^{2}}{\delta}+\sum_{t\in[T]\setminus\mathcal{E}_{T}}\left\{\mathbb{I}(a_{t}=\widehat{a}_{t})\left(\mathbb{E}_{t-1}[y_{\star,t}]-\mathbb{E}_{t-1}[y_{a_{t},t}]\right)\right\}\\ &\quad+\sum_{t\in[T]\setminus\mathcal{E}_{T}}\left\{\mathbb{I}(a_{t}\neq\widehat{a}_{t})\left(\mathbb{E}_{t-1}[y_{\star,t}]-\mathbb{E}_{t-1}[y_{a_{t},t}]\right)\right\}.\end{aligned}$ (30)
We first consider the second term in Eq.˜30. On the event $\{a_{t}=\widehat{a}_{t}\}$,
$\mathbb{E}_{t-1}[y_{\star,t}]-\mathbb{E}_{t-1}[y_{a_{t},t}]$ $=\widetilde{\mathbf{x}}_{a_{\star}}^{\top}{\boldsymbol{\mu}}_{\star}-\widetilde{\mathbf{x}}_{\widehat{a}_{t}}^{\top}{\boldsymbol{\mu}}_{\star}$
$=\widetilde{\mathbf{x}}_{a_{\star}}^{\top}\boldsymbol{\mu}_{\star}+\widetilde{\mathbf{x}}_{a_{\star}}^{\top}\widehat{\boldsymbol{\mu}}_{t-1}^{R}-\widetilde{\mathbf{x}}_{a_{\star}}^{\top}\widehat{\boldsymbol{\mu}}_{t-1}^{R}+\widetilde{\mathbf{x}}_{\widehat{a}_{t}}^{\top}\widehat{\boldsymbol{\mu}}_{t-1}^{R}-\widetilde{\mathbf{x}}_{\widehat{a}_{t}}^{\top}\widehat{\boldsymbol{\mu}}_{t-1}^{R}-\widetilde{\mathbf{x}}_{\widehat{a}_{t}}^{\top}{\boldsymbol{\mu}}_{\star}$
$\leq\left|\widetilde{\mathbf{x}}_{a_{\star}}^{\top}(\boldsymbol{\mu}_{\star}-\widehat{\boldsymbol{\mu}}_{t-1}^{R})\right|+\left|{\widetilde{\mathbf{x}}_{\widehat{a}_{t}}^{\top}(\boldsymbol{\mu}_{\star}-\widehat{\boldsymbol{\mu}}_{t-1}^{R})}\right|+\widetilde{\mathbf{x}}_{a_{\star}}^{\top}\widehat{\boldsymbol{\mu}}_{t-1}^{R}-\widetilde{\mathbf{x}}_{\widehat{a}_{t}}^{\top}\widehat{\boldsymbol{\mu}}_{t-1}^{R}$
$\leq 2\max_{a\in[K]}\left|\widetilde{\mathbf{x}}_{a}^{\top}\left({\boldsymbol{\mu}}_{\star}-\widehat{\boldsymbol{\mu}}_{t-1}^{R}\right)\right|+\widetilde{\mathbf{x}}_{a_{\star}}^{\top}\widehat{\boldsymbol{\mu}}_{t-1}^{R}-\widetilde{\mathbf{x}}_{\widehat{a}_{t}}^{\top}\widehat{\boldsymbol{\mu}}_{t-1}^{R}$
@@
$\leq\frac{4}{\sqrt{t}}\left(\frac{\sigma}{p}\sqrt{K\log\frac{t}{\delta}}+\sqrt{K}\right),$
with probability at least $1-3\delta$, for all $t$ such that $t\geq|\mathcal{E}_{t}|$ (Theorem˜4). Summing over $t$ gives, with probability at least $1-3\delta$,
$\sum_{t\in[T]\setminus\mathcal{E}_{T}}\left\{\mathbb{I}(a_{t}=\widehat{a}_{t})\left(\mathbb{E}_{t-1}[y_{\star,t}]-\mathbb{E}_{t-1}[y_{a_{t},t}]\right)\right\}\leq 8\sqrt{KT}\left(\frac{\sigma}{p}\sqrt{\log\frac{T}{\delta}}+1\right).$ (31)
For the last term in Eq.˜30, we have that
$\sum_{t\in[T]\setminus\mathcal{E}_{T}}\left\{\mathbb{I}(a_{t}\neq\widehat{a}_{t})\left(\mathbb{E}_{t-1}[y_{\star,t}]-\mathbb{E}_{t-1}[y_{a_{t},t}]\right)\right\}$
$\leq 2\left[\sum_{t\in[T]}\mathbb{I}(a_{t}\neq\widehat{a}_{t})-\mathbb{P}\left(a_{t}\neq\widehat{a}_{t}\right)+\mathbb{P}\left(a_{t}\neq\widehat{a}_{t}\right)\right]$
$\leq 2\sqrt{2T\log\frac{2}{\delta}}+\frac{4\sqrt{T}}{K-1}+4\delta,$ (32)
where the first term in Eq.˜32 holds with probability at least $1-\delta$ by Hoeffding’s inequality, and the remaining terms follow from Lemma˜5. Finally, by taking a union bound over Eq.˜31 and Eq.˜32, we obtain, with probability at least $1-4\delta$,
$\mathrm{Reg}(T)\leq\frac{32K^{2}}{(1-p)^{2}}\log\frac{2dT^{2}}{\delta}+2\sqrt{2T\log\frac{2}{\delta}}+\frac{4\sqrt{T}}{K-1}+4\delta+8\sqrt{KT}\left(\frac{\sigma}{p}\sqrt{\log\frac{T}{\delta}}+1\right),$